R HE = 20 sinf f2 kHz 0.47 kO 0.1 PUF Fig. 3.1

Consider the network
1. Determine the Zt
2. Find I
3. Calculate VR and VC.
4. Find P and Power Factor
Answer #1

R C 2 wc i VRCP) = (V. De Sinct Vect) - (V) sinkut-1/2) V=Vosinstad! (idzV), Cid Valle, no vo where; X = wc a fafe lor Peak c2= ceafeje te - Lets calculate Xe first : - 1 = 0.79617x10? Xc= FC 2%3:14X2Xlo?x0.1X166 Xc = 0. 79617X10% = 796.171 and R = 0VRCH = (OR) Sin (yoount+59.440) VRCH) = (2163810-PA) @47X1032) Sin (40001t+59.449) Ve(t) = 10-17 Sin (yooort+ 59.440) (Ve(u -